sum of ages of son and his father is 35 and the product is 150 find the ages
Answer by Cromlix(4381)

Hi there,
sum of ages of son and his father is 35.
Make son's age = x
Make father's age = y
x + y = 35........Eq(1)
the product is 150.
xy = 150..........Eq(2)
x = 150/y ........Eq(2)
Substitute 150/y from Eq(2) for x in Eq(1)
x + y = 35....Eq(1)
150/y + y = 35
Common denominator = y
150 + y^2 = 35y
Rearrange into trinomial
y^2 - 35y + 150 = 0
Factorize
(y - 30)(y - 5) = 0
y - 30 = 0
y = 30
(Disregard y - 5 = 0 => y = 5 as Father will be older than son)
Substitute y = 30 into Eq(1)
x + y = 35
x + 30 = 35
x = 35 - 30
x = 5
Boy's age = 5 years
Father's age = 30 years.
Hope this helps :-)